OpenStudy (anonymous):

Evaluate -15 + (-4).

OpenStudy (anonymous):

19?

OpenStudy (anonymous):

thank you

OpenStudy (chaotic_butterflies):

Remember, when adding a negative to another negative, just add the two integers together as you normally would if they were positive. The only difference is that in the end, it will always be negative. So, your answer is -19. :)

OpenStudy (anonymous):

negative plus a negative is a positive

OpenStudy (chaotic_butterflies):

@flutterSchoolSmokeWeed Remember there is nothing to cancel out the negative. A negative multiplied by a negative is positive, but it remains negative when adding a negative to another negative.

OpenStudy (chaotic_butterflies):

Don't you guys have calculators on your computers..? You can insert a negative number on it. And Openstudy automatically censors adult language.

OpenStudy (anonymous):

my calc put negatives in the wrong place and woooww, technology nowa days

OpenStudy (chaotic_butterflies):

Perhaps you inserted it incorrectly, it's not common for a calculator to malfunction unless it's older.

OpenStudy (anonymous):

(-2)*(-2) = -4 (-2)*(+2) = -4 (-4)+(-4)=-8 (-4 )+ (+8)= +4

OpenStudy (chaotic_butterflies):

Actually... (-2)*(-2)=4 (-2)*(2)= -4 (-4)+(-4)=-8 (-4 )+ (+8)= +4

OpenStudy (chaotic_butterflies):

All of yours were correct except the first :) @chathuranga